OpenStudy (anonymous):

Factor this x^2 – 8x + 15

OpenStudy (nathan917):

x^2-8x+15 = x^2-3x-5x+15 = x(x-3)-5(x-3) = (x-3)(x-5)

OpenStudy (anonymous):

a^2 – a – 20

OpenStudy (nathan917):

(a-5) (a+4)